# ADBM: Adversarial Diffusion Bridge Model for Reliable Adversarial Purification

### v1

@@ Abstract
Recently Diffusion-based Purification (DiffPure) has been recognized as an effective defense method against adversarial examples. However, we find DiffPure which directly employs the original pre-trained diffusion models for adversarial purification, to be suboptimal. This is due to an inherent trade-off between noise purification performance and data recovery quality. Additionally, the reliability of existing evaluations for DiffPure is questionable, as they rely on weak adaptive attacks. In this work, we propose a novel Adversarial Diffusion Bridge Model, termed ADBM. ADBM directly constructs a reverse bridge from the diffused adversarial data back to its original clean examples, enhancing the purification capabilities of the original diffusion models. Through theoretical analysis and experimental validation across various scenarios, ADBM has proven to be a superior and robust defense mechanism, offering significant promise for practical applications. Code will be made public soon.
@@ 1 Introduction
An intriguing problem in machine learning models, particularly Deep Neural Networks (DNNs), is the existence of adversarial examples [1, 2]. These examples introduce imperceptible adversarial perturbations leading to significant errors, which has posed severe threats to practical applications [3, 4]. Numerous methods have been proposed to defend against adversarial examples. But attackers can still evade most early methods by employing adaptive attacks [5, 6]. Adversarial Training (AT) methods [7, 8, 9, 10] are recognized as effective defense methods against adaptive attacks. However, AT typically involves re-training the entire DNNs using adversarial examples, which is impractical for real-world applications. Moreover, the effectiveness of AT is often limited to the specific attacks it has been trained against, making it brittle against unseen threats [11, 12].
@@
Recently, Adversarial Purification (AP) methods [13, 14] have gained increasing attention as they offer a potential solution to defend against unseen threats in a plug-and-play manner without retraining the classifiers. These methods utilize the so-called purification module, which exploits techniques such as generative models, as a pre-processing step to restore clean examples from adversarial examples, as illustrated in Figure 1(a). Recently, diffusion models [15], one type of generative model renowned for their efficacy, have emerged as potential AP solutions [16]. Diffusion models learn transformations between complex data distributions and simple distributions like the Gaussian distribution through forward diffusion and reverse prediction processes. In the context of adversarial defense, Diffusion-based Purification (DiffPure) [16], which tries to purify the adversarial examples by first adding Gaussian noise through the forward process with a small diffusion timestep and then recovering clean examples (removing the added Gaussian noise together with adversarial noise) through the reverse process, has achieved superior performance among recent AP methods.
However, we find that utilizing DiffPure in the context of AP is suboptimal. The primary reason is that DiffPure directly employs pre-trained diffusion models originally designed for generative tasks, rather than specifically for AP tasks. To delve deeper, the divergence in data distributions between clean and adversarial examples [17, 18] often leads to noticeable differences in their respective diffused distributions, while DiffPure relies on the assumption that the two diffused distributions are sufficiently close, such that the original reverse process can recover the diffused adversarial data distribution, as depicted in Figure 1(b). The conflict between the assumption and the actual situation compromises the effectiveness of AP. DiffPure attempts to mitigate this conflict by introducing significant noise with a large diffusion timestep. Nevertheless, this solution is impractical; a substantial diffusion timestep can severely corrupt the global structure of the input, resulting in a fundamental trade-off between noise purification efficiency and recovery quality. In addition, we observe that existing evaluations for DiffPure rely on weak adaptive attacks, which may inadvertently give a false sense of security regarding diffusion-based purification.
To address the aforementioned problems, this study performs a systematic investigation. We first establish a reliable adaptive attack evaluation method for diffusion-based purification. With the rigorous evaluation, our preliminary findings suggest that the robustness of DiffPure is overestimated in existing works [16, 19, 20]. To improve the robustness of diffusion-based purification, we then propose a novel Adversarial Diffusion Bridge Model, termed ADBM. Unlike original diffusion models relying on the similarity between the diffused distributions of clean and adversarial examples for a balanced trade-off, ADBM constructs a direct reverse process (or “bridge”) from the diffused adversarial data distribution to the distribution of clean examples, as shown in Figure 1(c). The theoretical analysis supports the superiority of ADBM over using original pre-trained diffusion models (DiffPure). In addition, we discuss how to accelerate ADBM for efficient adversarial purification to enhance the practicality of ADBM as a defense mechanism.
Experimental results show that ADBM achieved better adversarial robustness than DiffPure under reliable adaptive attacks. In particular, ADBM achieved a 4.4% robustness gain compared with DiffPure on average on CIFAR-10 [21], while the clean accuracies kept comparable. ADBM also demonstrated competitive performances with AT methods for seen adversarial threats and stood out among recent AP methods. Furthermore, ADBM exhibited much better adversarial robustness than AT methods when facing unseen threats. Additionally, the evaluation results against transfer-based and query-based attacks indicate the practicality of ADBM compared with existing methods.
@@
  • Experiments in various attack settings validate the effectiveness of ADBM, highlighting its reliability and potential as a defense method for practical scenarios when compared with AT.
@@ 2 Preliminary and Related Work
Diffusion Models. Given a data distribution $q({\mathbf{x}}_{0})$, DDPM [15] constructs a discrete-time Markov chain $\{{\mathbf{x}}_{0},\ldots,{\mathbf{x}}_{T}\}$ as the forward process for ${\mathbf{x}}_{0}\sim q({\mathbf{x}}_{0})$. Gaussian noise is gradually added to ${\mathbf{x}}_{0}$ during the forward process following a scaling schedule $\{\beta_{0},\beta_{1},\cdots,\beta_{T}\}$, where $\beta_{0}=0$ and $\beta_{T}\rightarrow 1$, such that ${\mathbf{x}}_{T}$ is near an isotropic Gaussian distribution:
@@
$q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ $={\mathcal{N}}({\mathbf{x}}_{t};\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0},(1-\bar{\alpha}_{t}){\mathbf{I}}),\textit{i.e.},{\mathbf{x}}_{t}({\mathbf{x}}_{0},{\boldsymbol{\epsilon}})=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},{\boldsymbol{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\mathbf{I}}).$ (2)
@@
${\mathbf{x}}_{\theta}({\mathbf{x}}_{t},t):=({\mathbf{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\mathbf{x}}_{t},t))/\sqrt{\bar{\alpha}_{t}}.$ (3)
@@
$L_{d}$ $=\mathbb{E}_{{\boldsymbol{\epsilon}},t}\left[\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},t)\|^{2}\right].$ (4)
To accelerate the reverse process of DDPM, which typically involves hundreds of steps, Song et al. [22] propose a DDIM sampler based on the intuition that the multiple reverse steps can be performed at a single step via a non-Markov process. Song et al. [23] generalize the discrete-time diffusion model to continuous-time from the Stochastic Differential Equation (SDE) perspective.
Diffusion Models for Adversarial Robustness. Recent studies have demonstrated the efficacy of diffusion models [15] in enhancing adversarial robustness in several ways. Some researches leverage much data generated by diffusion models to improve the AT performance [24, 25], but these AT-based methods do not generalize well under unseen threat models. Chen et al. [20] show that a single diffusion model can be transformed into an adversarially robust classifier using Bayes’ rule, but at thousands of times the inference cost. DiffPure [16] employs a diffusion model as a plug-and-play pre-processing module to purify adversarial noise. Wang et al. [26] improved DiffPure by employing inputs to guide the reverse process of the diffusion model to ensure the purified examples are close to input examples. Zhang et al. [27] improved DiffPure by incorporating the reverse SDE with multiple Langevin dynamic runs. Zhang et al. [28] maximized the evidence lower bound of the likelihood estimated by diffusion models to increase the likelihood of corrupted images. DiffPure has also shown potential in improving certified robustness within the framework of randomized smoothing [29, 30]. Nevertheless, the practicality of randomized smoothing is greatly hindered by the time-consuming Monte Carlo sampling [31]. Different from these works, we present an diffusion-based purification method for empirical robustness in practical scenarios.
@@ 3 Reliable Evaluation for DiffPure
Before delving into the details of ADBM, it is important to discuss the white-box adaptive attack for diffusion-based purification first. As the original implementation of DiffPure needs dozens of prediction steps in the reverse process, it is challenging to compute the full gradient of the whole purification process due to memory constraints. To evaluate the robustness of diffusion-based purification, several techniques have been proposed. But we found that the adaptive evaluations for DiffPure remained insufficient, as detailed in Sec. A.1. We build on these previous insights [16, 20, 19] to develop a straightforward yet effective adaptive attack method against diffusion-based purification. We employ the gradient-based PGD attack [7], utilizing the full gradient calculation via gradient-checkpointing and incorporating a substantial number of EOT [5] and iteration steps. Especially, we set the PGD iteration steps to 200, with 20 EOT samples for each iteration. We note that the cost of reliable evaluation is high yet worthwhile to avoid a false sense of security on DiffPure, as discussed in Sec. A.2. Tab. 2 shows the results of DiffPure with a DDPM sampler under different evaluations. [^1] The DDPM sampler represents one discretization version of the reverse VP-SDE sampler [23], which is originally employed in DiffPure. Their differences become negligible at a large timestep, allowing us to use VP-SDE and DDPM interchangeably throughout our work. Consistent with Nie et al. [16], we conducted the adaptive attack three times on a subset of 512 randomly sampled images from the test set of CIFAR-10. The results demonstrate that our attack significantly reduced the reported robustness of DiffPure, lowering it from 71.29% and 51.13% to 45.83% when compared with the originally reported results [16] and the recent best practice [19], respectively. We found that increasing the number of iterations or EOT steps further did not lead to higher attack success rates.
@@
Reverse steps. DiffPure suffers from a high inference cost, which greatly hinders its practical application. The reverse step directly impacts the inference cost of DiffPure. Following Lee and Kim [19], we evaluated both the original DDPM sampler and the DDIM sampler. We present the results in Tab. 2 using different reverse steps and samplers (standard deviation omitted, about 1%). The results show that decreasing the number of reverse steps significantly reduced both clean accuracy and robustness when using DDPM. However, for the DDIM sampler, although clean accuracy slightly decreased with fewer reverse steps, robustness was not significantly affected. This can be attributed to the fact that DDIM does not introduce additional randomness during the reverse process.
Forward steps. We observed a continuous decrease in clean accuracy and a continuous increase in robustness as the number of forward steps increased. Details can be found in Sec. A.3.
@@ 4.1 Training Objective
ADBM is a diffusion model specifically designed for purifying adversarial noise. It adopts a forward process similar to DDPM, with the difference that ADBM assumes the existence of adversarial noise ${\boldsymbol{\epsilon}}_{a}$ at the starting point of the forward process during training. This means that the starting point of the forward process is ${\mathbf{x}}_{0}^{a}={\mathbf{x}}_{0}+{\boldsymbol{\epsilon}}_{a}$ for each ${\mathbf{x}}_{0}$. Thus, according to Eq. 2, the forward process can be represented as ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},{\boldsymbol{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\mathbf{I}}),0\leq t\leq T$, where $T$ denotes the actual forward timestep when performing AP. $T$ is typically set to a lower value for AP, e.g., 100 in DiffPure, than that used in generative tasks, e.g., 1,000, to avoid completely corrupting ${\mathbf{x}}_{0}$. We discuss how to obtain ${\boldsymbol{\epsilon}}_{a}$ in Sec. 4.2.
In the reverse process of ADBM, the objective is to learn a Markov chain $\{\hat{{\mathbf{x}}}_{t}\}_{t:T\to 0}$ that can directly transform from the diffused adversarial data distribution (i.e., ${\mathbf{x}}^{a}_{T}$) to the clean data distribution (i.e., ${\mathbf{x}}_{0}$), as shown in Figure 1(c). To achieve this, the starting point and the end point of $\hat{{\mathbf{x}}}_{t}$ should be defined as $\hat{{\mathbf{x}}}_{T}:={\mathbf{x}}^{a}_{T}$ and $\hat{{\mathbf{x}}}_{0}:={\mathbf{x}}_{0}$, respectively. Notably, $\hat{{\mathbf{x}}}_{T}$ contains adversarial noise, while $\hat{{\mathbf{x}}}_{0}$ does not. To explicitly align the trajectory of $\hat{{\mathbf{x}}}_{t}$ with the starting and ending points, we introduce a coefficient $k_{t}$, such that $\hat{{\mathbf{x}}}_{t}:={\mathbf{x}}_{t}^{d}={\mathbf{x}}_{t}^{a}-k_{t}{\boldsymbol{\epsilon}}_{a}$ for $0\leq t\leq T$, with $k_{0}=1$ and $k_{T}=0$. With Bayes’ rule and the property of Gaussian distribution, we have
@@
where $A$ is a constant independent of ${\boldsymbol{\epsilon}}_{a}$, $B$ is the coefficient of ${\mathbf{x}}_{t-1}^{d}$ dependent on ${\boldsymbol{\epsilon}}_{a}$, and $C$ is a term without ${\mathbf{x}}_{t-1}^{d}$. In inference, we expect that ${\boldsymbol{\epsilon}}_{a}$ in Eq. 5 can be eliminated since only ${\mathbf{x}}_{0}^{a}$ is given by attackers and ${\boldsymbol{\epsilon}}_{a}$ cannot be decoupled from ${\mathbf{x}}_{0}^{a}$ directly. Based on the property of Gaussian distribution, eliminating all terms related to ${\boldsymbol{\epsilon}}_{a}$ in $B$ and $C$ in Eq. 5 can be achieved by eliminating all terms related to ${\boldsymbol{\epsilon}}_{a}$ in $B$. This yields:
$\frac{\sqrt{\alpha_{t}}(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\boldsymbol{\epsilon}}_{a}}{1-\alpha_{t}}-\frac{(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\boldsymbol{\epsilon}}_{a}}{1-\bar{\alpha}_{t-1}}=0.$ (6)

### v3

@@ Abstract
Recently Diffusion-based Purification (DiffPure) has been recognized as an effective defense method against adversarial examples. However, we find DiffPure which directly employs the original pre-trained diffusion models for adversarial purification, to be suboptimal. This is due to an inherent trade-off between noise purification performance and data recovery quality. Additionally, the reliability of existing evaluations for DiffPure is questionable, as they rely on weak adaptive attacks. In this work, we propose a novel Adversarial Diffusion Bridge Model, termed ADBM. ADBM directly constructs a reverse bridge from the diffused adversarial data back to its original clean examples, enhancing the purification capabilities of the original diffusion models. Through theoretical analysis and experimental validation across various scenarios, ADBM has proven to be a superior and robust defense mechanism, offering significant promise for practical applications. Code is available at https://github.com/LixiaoTHU/ADBM.
@@ 1 Introduction
An intriguing problem in machine learning models, particularly Deep Neural Networks (DNNs), is the existence of adversarial examples (Szegedy et al., 2014; Goodfellow et al., 2015). These examples introduce imperceptible adversarial perturbations leading to significant errors, which has posed severe threats to practical applications (Eykholt et al., 2018; Li et al., 2025). Numerous methods have been proposed to defend against adversarial examples. But attackers can still evade most early methods by employing adaptive attacks (Athalye et al., 2018; Tramèr et al., 2020). Adversarial Training (AT) methods (Madry et al., 2018; Li et al., 2024c; b) are recognized as effective defense methods against adaptive attacks. However, AT typically involves re-training the entire DNNs using adversarial examples, which is impractical for real-world applications. Moreover, the effectiveness of AT is often limited to the specific attacks it has been trained against, making it brittle against unseen threats (Zhang et al., 2019; Laidlaw et al., 2021).
Recently, Adversarial Purification (AP) methods (Shi et al., 2021; Yoon et al., 2021) have gained increasing attention as they offer a potential solution to defend against unseen threats in a plug-and-play manner without retraining the classifiers. These methods utilize the so-called purification module, which exploits techniques such as generative models, as a pre-processing step to restore clean examples from adversarial examples, as illustrated in Figure 1(a). Recently, diffusion models (Ho et al., 2020), one type of generative model renowned for their efficacy, have emerged as potential AP solutions (Nie et al., 2022). Diffusion models learn transformations between complex data distributions and simple distributions like the Gaussian distribution through forward diffusion and reverse prediction processes. In the context of adversarial defense, Diffusion-based Purification (DiffPure) (Nie et al., 2022), which tries to purify the adversarial examples by first adding Gaussian noise through the forward process with a small diffusion timestep and then recovering clean examples (removing the added Gaussian noise together with adversarial noise) through the reverse process, has achieved superior performance among recent AP methods.
However, we find that utilizing DiffPure in the context of AP is suboptimal. The primary reason is that DiffPure directly employs pre-trained diffusion models originally designed for generative tasks, rather than specifically for AP tasks. To delve deeper, the divergence in data distributions between clean and adversarial examples (Stutz et al., 2019; Xie et al., 2020) often leads to noticeable differences in their respective diffused distributions, while DiffPure relies on the assumption that the two diffused distributions are sufficiently close, such that the original reverse process can recover the diffused adversarial data distribution, as depicted in Figure 1(b). The conflict between the assumption and the actual situation compromises the effectiveness of AP. DiffPure attempts to mitigate this conflict by introducing significant noise with a large diffusion timestep. Nevertheless, this solution is impractical; a substantial diffusion timestep can severely corrupt the global structure of the input, resulting in a fundamental trade-off between noise purification efficiency and recovery quality. In addition, we observe that existing evaluations for DiffPure rely on weak adaptive attacks, which may inadvertently give a false sense of security regarding diffusion-based purification.
@@
To address the aforementioned problems, this study performs a systematic investigation. We first establish a reliable adaptive attack evaluation method for diffusion-based purification. With the rigorous evaluation, our preliminary findings suggest that the robustness of DiffPure is overestimated in existing works (Nie et al., 2022; Lee & Kim, 2023; Chen et al., 2023a). To improve the robustness of diffusion-based purification, we then propose a novel Adversarial Diffusion Bridge Model, termed ADBM. Unlike original diffusion models relying on the similarity between the diffused distributions of clean and adversarial examples for a balanced trade-off, ADBM constructs a direct reverse process (or “bridge”) from the diffused adversarial data distribution to the distribution of clean examples, as shown in Figure 1(c). The theoretical analysis supports the superiority of ADBM over using original pre-trained diffusion models (DiffPure). In addition, we discuss how to accelerate ADBM for efficient adversarial purification to enhance the practicality of ADBM as a defense mechanism.
Experimental results show that ADBM achieved better robustness than DiffPure under reliable adaptive attacks. In particular, ADBM achieved a 4.4% robustness gain compared with DiffPure on average on CIFAR-10 (Krizhevsky et al., 2009), while the clean accuracies kept comparable. ADBM also demonstrated competitive performances with AT methods for seen adversarial threats and stood out among recent AP methods. Furthermore, ADBM exhibited much better adversarial robustness than AT methods when facing unseen threats. Additionally, the evaluation results against transfer-based and query-based attacks indicate the practicality of ADBM compared with existing methods.
@@
  • Experiments in various settings validate the effectiveness of ADBM, highlighting its reliability and potential as a defense method for practical scenarios when compared with AT.
@@ 2 Preliminary and Related Work
Diffusion models. Given a data distribution $q({\mathbf{x}}_{0})$, DDPM (Ho et al., 2020) constructs a discrete-time Markov chain $\{{\mathbf{x}}_{0},\ldots,{\mathbf{x}}_{T}\}$ as the forward process for ${\mathbf{x}}_{0}\sim q({\mathbf{x}}_{0})$. Gaussian noise is gradually added to ${\mathbf{x}}_{0}$ during the forward process following a scaling schedule $\{\beta_{0},\beta_{1},\cdots,\beta_{T}\}$, where $\beta_{0}=0$ and $\beta_{T}\rightarrow 1$, such that ${\mathbf{x}}_{T}$ is near an isotropic Gaussian distribution:
@@
$q({\mathbf{x}}_{t}|{\mathbf{x}}_{0})$ $={\mathcal{N}}({\mathbf{x}}_{t};\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0},(1-\bar{\alpha}_{t}){\mathbf{I}}),\textit{i.e.},{\mathbf{x}}_{t}({\mathbf{x}}_{0},{\bm{\epsilon}})=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\mathbf{I}}).$ (2)
@@
${\mathbf{x}}_{\theta}({\mathbf{x}}_{t},t):=({\mathbf{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}}_{\theta}({\mathbf{x}}_{t},t))/\sqrt{\bar{\alpha}_{t}}.$ (3)
@@
$L_{d}$ $=\mathbb{E}_{{\bm{\epsilon}},t}\left[\|{\bm{\epsilon}}-{\bm{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},t)\|^{2}\right].$ (4)
To accelerate the reverse process of DDPM, which typically involves hundreds of steps, Song et al. (2020) propose a DDIM sampler based on the intuition that the multiple reverse steps can be performed at a single step via a non-Markov process. Song et al. (2021) generalize the discrete-time diffusion model to continuous-time from the Stochastic Differential Equation (SDE) perspective.
Another series of works seemly related to ours are diffusion bridges (Zhou et al., 2024; Zheng et al., 2024; Bortoli et al., 2021; Liu et al., 2023b), which enlarge the design space of diffusion models by interpolating between two paired distributions. However, ADBM and current diffusion bridge models have distinct motivations and designs. Current diffusion bridge models either map the data distribution to a simple prior distribution (Bortoli et al., 2021) or directly create a Gaussian bridge (Zhou et al., 2024) or an optimal transport bridge (Zheng et al., 2024; Liu et al., 2023b) between two data distributions. The former approach (Bortoli et al., 2021) fundamentally cannot be used for AP, while the latter methods (Zhou et al., 2024; Bortoli et al., 2021; Liu et al., 2023b) tend to a theoretical training collapse for the AP task. We give a detailed discussion to these works in Sec. A.1.
Diffusion models for adversarial robustness. Recent studies have demonstrated the efficacy of diffusion models (Ho et al., 2020) in enhancing adversarial robustness in several ways. Some researches leverage much data generated by diffusion models to improve the AT performance (Gowal et al., 2021; Wang et al., 2023), but these AT-based methods do not generalize well under unseen threat models. DiffPure (Nie et al., 2022) employs a diffusion model as a plug-and-play pre-processing module to purify adversarial noise. Wang et al. (2022) improved DiffPure by employing inputs to guide the reverse process of the diffusion model to ensure the purified examples are close to input examples. Zhang et al. (2023) improved DiffPure by incorporating the reverse SDE with multiple Langevin dynamic runs. Zhang et al. (2024) maximized the evidence lower bound of the likelihood estimated by diffusion models to increase the likelihood of corrupted images. DiffPure has also shown potential in improving certified robustness within the framework of randomized smoothing (Carlini et al., 2023; Xiao et al., 2023). Nevertheless, the practicality of randomized smoothing is greatly hindered by the time-consuming Monte Carlo sampling (Cohen et al., 2019). Different from these works, we present an diffusion-based purification method for empirical robustness in practical scenarios. In addition to these works, Chen et al. (2023a) show that a single diffusion model can be transformed into an adversarially robust diffusion classifier (RDC) using Bayes’ rule. However, compared with diffusion-based purification, RDC requires thousands of times the inference cost and may not scale up, limiting its practical usefulness. We discuss it further in Sec. A.2.
@@ 3 Reliable Evaluation for DiffPure
Before delving into the details of ADBM, it is important to discuss the white-box adaptive attack for diffusion-based purification first. As the original implementation of DiffPure needs dozens of prediction steps in the reverse process, it is challenging to compute the full gradient of the whole purification process due to memory constraints. To evaluate the robustness of diffusion-based purification, several techniques have been proposed. But we found that the adaptive evaluations for DiffPure remained insufficient, as detailed in Sec. B.1. We build on these previous insights (Nie et al., 2022; Chen et al., 2023a; Lee & Kim, 2023) to develop a straightforward yet effective adaptive attack method against diffusion-based purification. We employ the gradient-based PGD attack (Madry et al., 2018), utilizing the full gradient calculation via gradient-checkpointing and incorporating a substantial number of EOT (Athalye et al., 2018) and iteration steps. Especially, we set the PGD iteration steps to 200, with 20 EOT samples for each iteration. We note that the cost of reliable evaluation is high yet worthwhile to avoid a false sense of security on DiffPure, as discussed in Sec. B.2. Tab. 2 shows the results of DiffPure with a DDPM sampler under different evaluations. [^1] The DDPM sampler represents one discretization version of the reverse VP-SDE sampler (Song et al., 2021), which is originally employed in DiffPure. Their differences become negligible at a large timestep, allowing us to use VP-SDE and DDPM interchangeably throughout our work. Consistent with Nie et al. (2022), we conducted the adaptive attack three times on a subset of 512 randomly sampled images from the test set of CIFAR-10. The results demonstrate that our attack significantly reduced the reported robustness of DiffPure, lowering it from 71.29% and 51.13% to 45.83% when compared with the originally reported results (Nie et al., 2022) and the recent best practice (Lee & Kim, 2023), respectively. We found that increasing the number of iterations or EOT steps further did not lead to higher attack success rates.
@@
Reverse steps. DiffPure suffers from a high inference cost, which greatly hinders its practical application. The reverse step directly impacts the inference cost of DiffPure. Following Lee & Kim (2023), we evaluated both the original DDPM sampler and the DDIM sampler. We present the results in Tab. 2 using different reverse steps and samplers (standard deviation omitted, about 1%). The results show that decreasing the number of reverse steps significantly reduced both clean accuracy and robustness when using DDPM. However, for the DDIM sampler, although clean accuracy slightly decreased with fewer reverse steps, robustness was not significantly affected. This can be attributed to the fact that DDIM does not introduce additional randomness during the reverse process.
Forward steps. We observed a continuous decrease in clean accuracy and a continuous increase in robustness as the number of forward steps increased. Details can be found in Sec. B.3.
@@ 4.1 Training Objective
ADBM is a diffusion model specifically designed for purifying adversarial noise. It adopts a forward process similar to DDPM, with the difference that ADBM assumes the existence of adversarial noise ${\bm{\epsilon}}_{a}$ at the starting point of the forward process during training. This means that the starting point of the forward process is ${\mathbf{x}}_{0}^{a}={\mathbf{x}}_{0}+{\bm{\epsilon}}_{a}$ for each ${\mathbf{x}}_{0}$. Thus, according to Eq. 2, the forward process can be represented as ${\mathbf{x}}_{t}^{a}=\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}^{a}+\sqrt{1-\bar{\alpha}_{t}}{\bm{\epsilon}},{\bm{\epsilon}}\sim{\mathcal{N}}({\bm{0}},{\mathbf{I}}),0\leq t\leq T$, where $T$ denotes the actual forward timestep when performing AP. $T$ is typically set to a lower value for AP, e.g., 100 in DiffPure, than that used in generative tasks, e.g., 1,000, to avoid completely corrupting ${\mathbf{x}}_{0}$. We discuss how to obtain ${\bm{\epsilon}}_{a}$ in Sec. 4.2.
In the reverse process of ADBM, the objective is to learn a Markov chain $\{\hat{{\mathbf{x}}}_{t}\}_{t:T\to 0}$ that can directly transform from the diffused adversarial data distribution (i.e., ${\mathbf{x}}^{a}_{T}$) to the clean data distribution (i.e., ${\mathbf{x}}_{0}$), as shown in Figure 1(c). To achieve the goal of purification, we expect that the starting point and the end point of $\hat{{\mathbf{x}}}_{t}$ to be $\hat{{\mathbf{x}}}_{T}:={\mathbf{x}}^{a}_{T}$ and $\hat{{\mathbf{x}}}_{0}:={\mathbf{x}}_{0}$, respectively. Notably, $\hat{{\mathbf{x}}}_{T}$ contains adversarial noise, while $\hat{{\mathbf{x}}}_{0}$ does not. To explicitly align the trajectory of $\hat{{\mathbf{x}}}_{t}$ with the starting and ending points, we introduce a coefficient $k_{t}$, expecting that $\hat{{\mathbf{x}}}_{t}:={\mathbf{x}}_{t}^{d}={\mathbf{x}}_{t}^{a}-k_{t}{\bm{\epsilon}}_{a}$ for $0\leq t\leq T$, with $k_{0}=1$ and $k_{T}=0$. With Bayes’ rule and the property of Gaussian distribution, we have
@@
where $A$ is a constant independent of ${\bm{\epsilon}}_{a}$, $B$ is the coefficient of ${\mathbf{x}}_{t-1}^{d}$ dependent on ${\bm{\epsilon}}_{a}$, and $C$ is a term without ${\mathbf{x}}_{t-1}^{d}$. In inference, we expect that ${\bm{\epsilon}}_{a}$ in Eq. 5 can be eliminated since only ${\mathbf{x}}_{0}^{a}$ is given by attackers and ${\bm{\epsilon}}_{a}$ cannot be decoupled from ${\mathbf{x}}_{0}^{a}$ directly. Based on the property of Gaussian distribution, eliminating all terms related to ${\bm{\epsilon}}_{a}$ in $B$ and $C$ in Eq. 5 can be achieved by eliminating all terms related to ${\bm{\epsilon}}_{a}$ in $B$. This yields:
$\frac{\sqrt{\alpha_{t}}(\sqrt{\alpha_{t}}k_{t-1}-k_{t}){\bm{\epsilon}}_{a}}{1-\alpha_{t}}-\frac{(\sqrt{\bar{\alpha}_{t-1}}-k_{t-1}){\bm{\epsilon}}_{a}}{1-\bar{\alpha}_{t-1}}=0.$ (6)